OpenStudy (anonymous):

f(h)=-6h-2/3,find f(-3)

OpenStudy (saifoo.khan):

52/3

OpenStudy (anonymous):

f(h)=-6h-2/3,find f(-3) f(-3)=-6(-3)-2/3=18-2/3=16/3

OpenStudy (anonymous):

18 - (2/3) (54/3) - (2/3) 52/3

OpenStudy (anonymous):

sorry my bad i thought whole was divide by three